40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 18 22:35:23 2019, gautam, Update, SUS, ETMY actuator diagnosis  
    Reply  Fri Apr 19 11:36:23 2019, gautam, Update, SUS, No consistent solution for output matrix 
       Reply  Fri Apr 19 15:13:38 2019, rana, Update, SUS, No consistent solution for output matrix 
          Reply  Fri Apr 19 16:19:42 2019, gautam, Update, SUS, Actuation matrix still not orthogonal 
             Reply  Fri Apr 19 19:22:15 2019, rana, Update, SUS, Actuation matrix still not orthogonal 
    Reply  Mon Apr 22 22:43:15 2019, gautam, Update, SUS, ETMY sensor diagnosis  ETMY_sensorSpectra_consolidated.pdf
       Reply  Thu Apr 25 00:30:45 2019, gautam, Update, SUS, ETMY BR mode  ETMY_sensorSpectra_BRmode.pdfETMY_sensorSpectra_BRmode.pdfETMX_sensorSpectra_BRmode.pdfIMG_5993.JPG
Message ID: 14562     Entry time: Mon Apr 22 22:43:15 2019     In reply to: 14551     Reply to this: 14569
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY sensor diagnosis  

Here are the results from this test. The data for 17 April is with the DC bias for ETMY set to the nominal values (which gives good Y arm cavity alignment), while on 18 April, I changed the bias values until all four shadow sensors reported values that were at least 100 cts different from 17 April. The times are indicated in the plot titles in case anyone wants to pull the data (I'll point to the directory where they are downloaded and stored later).

There are 3 visible peaks. There was negligible shift in position (<5 mHz)  / change in Q of any of these with the applied Bias voltage. I didn't attempt to do any fitting as it was not possible to determine which peak corresponds to which DoF by looking at the complex TFs between coils (at each peak, different combinations of 3 OSEMs have the same phase, while the fourth has ~180 deg phase lead/lag). FTR, the wiki leads me to expect the following locations for the various DoFs, and I've included the closest peak in the current measured data in parentheses:

DoF Frequency [Hz]
POS 0.982 (0.947)
PIT 0.86 (0.886)
YAW 0.894 (0.886)
SIDE 1.016 (0.996)

However, this particular SOS was re-suspended in 2016, and this elog reports substantially different peak positions, in particular, for the YAW DoF (there were still 4). The Qs of the peaks from last week's measurements are in the range 250-350.


Repeat the free-swinging ringdown with the ETMY bias voltage adjusted such that all the OSEM PDmons report ~100 um different position from the "nominal" position (i.e. when the Y arm cavity is aligned). Investigate whether the resulting eigenmode frequencies / Qs are radically different. I'm setting the optic free-swinging on my way out tonight. Optic kicked at 1239690286.

Attachment 1: ETMY_sensorSpectra_consolidated.pdf  539 kB  Uploaded Mon Apr 22 23:48:50 2019  | Hide | Hide all
ETMY_sensorSpectra_consolidated.pdf ETMY_sensorSpectra_consolidated.pdf
ELOG V3.1.3-